# Fast Feature Matching of UAV Images via Matrix Band Reduction-based GPU Data Schedule

### v1

@@ 3.4.2 SfM-based 3D reconstruction
Metric Method Dataset 1 Dataset2 Dataset 3 Time cost (min) ColMap-GPU 13.9 46.1 703.2 TLBDS 14.8 44.6 714.3 Metashape 41.8 69.9 Pix4Dmapper 234.2 352.4 Ours 13.9 52.4 669.5 #Registered images ColMap-GPU 3,736 4,029 21,624 TLBDS 3,734 4,027 21,349 Metashape 3,743 4,029 Pix4Dmapper 3736 4027 Ours 3,728 4,027 21,627 #3D points memory costs ColMap-GPU 874,869 1,340,548 7,698,043 TLBDS 898,650 1,394,779 8,063,822 Metashape 2,824,443 2,532,249 Pix4Dmapper 4,407,343 5,206,527 Ours 896,669 1,403,623 8,100,207
@@ 4 Conclusions
This study proposes a matrix band reduction-based GPU data schedule for fast feaure matching of UAV images. The core idea is to divide the whole dataset into blocks based on the matrix band reduction and achieve efficient feature matching via GPU-accelerated cascade hashing. First, by using the VLAD-HNSW image retrieval technique, a view graph is created to establish image connections, which is then divided into compact blocks via the matrix band reduction based data schedule strategy. The data schedule strategy adapts well to the connection structure of images and the volume of GPU memory. Second, guided by the generated blocks, feature matching is executed sequentially based on GPU-accelerated cascade hashing, in which initial candidate matches are refined by combining SAO-based local constraint and RANSAC-based global verification in outlier removal. Finally, the proposed algorithm is evaluated and compared by using large-scale UAV datasets. The experimental results demonstrate that the proposed solution can achieve efficient feature matching with tens to hundreds speedup ratios and provide reliable matches for SfM-based 3D reconstruction.
@@ Acknowledgments
This research was funded by the National Natural Science Foundation of China (Grant No. 42371442), and the Hubei Provincial Natural Science Foundation of China (Grant No. 2023AFB568).

### v2

@@ 3.5.1 Relative BA without GCPs
Metric Method Dataset 1 Dataset2 Dataset 3 Time cost (min) ColMap-GPU 13.9 46.1 703.2 TLBDS 14.8 44.6 714.3 Metashape 41.8 69.9 Pix4Dmapper 234.2 352.4 Ours 13.9 52.4 669.5 #Registered images ColMap-GPU 3,736 4,029 21,624 TLBDS 3,734 4,027 21,349 Metashape 3,743 4,029 Pix4Dmapper 3736 4027 Ours 3,728 4,027 21,627 #3D points ColMap-GPU 874,869 1,340,548 7,698,043 TLBDS 898,650 1,394,779 8,063,822 Metashape 2,824,443 2,532,249 Pix4Dmapper 4,407,343 5,206,527 Ours 896,669 1,403,623 8,100,207
@@
For relative BA without GCPs, SfM is used to compare the efficiency and completeness of reconstructed models, in which the efficiency is quantified by time costs, and the completeness is measured by the number of registered images and 3D points. In this test, the parallel SfM solution proposed in Jiang et al. (2024a) is used as the engine for ColMap-GPU, TLBDS, and the proposed algorithm.
@@ 3.5.2 Absolute BA with GCPs
4 Discussion
In this study, we propose a GPU data schedule algorithm for efficient feature matching of large-scale UAV images. The main issue addressed by the proposed algorithm is to make a balance between redundant data IO burden and the usage of GPU computing power. For this purpose, two major contributions have been made to the proposed solution. On the one hand, by using VLAD-HNSW image retrieval for view graph construction, we propose a MBR based data schedule strategy to divide the sparsely connected view graph into compact sub-graphs. The data schedule strategy adapts well to the volume of GPU memory as well as the connection structure of images; on the other hand, using the data schedule strategy, we further implement a cascade hashing based feature matching workflow, which integrates the SAO-based local constraint and RANSAC-based global verification for efficient and reliable outlier removal. By using large-scale UAV datasets captured by classical oblique and optimized views photogrammetry, the performance of the proposed algorithm has been extensively evaluated in terms of feature matching and SfM-based 3D reconstruction and compared with state-of-the-art open-source and commercial software packages.
For the proposed algorithm, there are two critical parameters, i.e., the schedule block size $Size_{blk}$ for data schedule and the ratio test threshold $t_{r}$ for outlier removal. According to the test results presented in Section 3.2, we can see that for UAV datasets with dense connections, the schedule block size influences the overall efficiency of feature matching, and a large value can increase the feature matching efficiency; on the contrary, it almost has no influence on the datasets with sparse connections, as verified by the results in Figure 9. In practice, a majority of UAV datasets have been captured according to the rule the classical oblique photogrammetry. The proposed algorithm can perform well with a proper schedule block size that fits the GPU memory volume. For the ratio test threshold, we can see that the optimal value of 0.5 has been verified by the results presented in Figure 9, which differs from the setting in other libraries, e.g., 0.8 in SIFTGPU. The main reason can be explained by the pre-filtering of cascade hashing mapping. In addition, the experimental results in Section 3.4 demonstrate that the proposed solution achieves the highest efficiency in feature matching and provides reliable matching results for SfM-based 3D reconstruction. On the one hand, it can obviously increase the efficiency of feature matching with speedup ratios ranging from 77.0 to 100.0 compared with traditional KD-Tree based matching methods. The major advantages lie in the MBR-based data schedule and GPU-accelerated cascade hashing; on the other hand, the proposed solution can achieve comparable relative BA results in terms of the number of registered images and 3D points and better performance in absolute BA when compared with ColMap-GPU and TLBDS. In conclusion, the proposed algorithm can be an efficient and reliable solution for feature matching of large-scale UAV images.
For the proposed feature matching algorithm, some issues should be paid attention. First, VLAD-HNSW based image retrieval has been designed for match pair selection, and its performance has been validated using large-scale UAV datasets. The sparsity of image networks has an impact on MBR-based feature matching performance, but not on image retrieval performance. As an extreme case, the linear flight path causes the image adjacency matrix with non-zero values only along the diagonal. In this case, the data schedule problem is simplified, which indicates that the MBR-based data schedule strategy has no obvious advantage over other algorithms. Second, feature matching based on cascade hashing would perform worse in the presence of low texture or repeating patterns. Almost every other feature matching algorithm does the same. The SAO constraint had been designed for this study in order to remove outliers following the initial feature matching. SAO-based outlier elimination can also be effective, even when low-texture or repeating patterns would reduce the quantity and inlier ratio of initial matches.
@@ 5 Conclusions
This study proposes matrix band reduction-based GPU data schedule for efficient feaure matching of UAV images. The core idea is to divide the whole dataset into blocks via matrix band reduction and achieve efficient feature matching based on GPU-accelerated cascade hashing. First, by using the VLAD-HNSW image retrieval technique, a view graph is created to establish image connections, which is then divided into compact blocks via matrix band reduction. The data schedule strategy adapts well to the connection structure of images and the volume of GPU memory. Second, guided by the generated blocks, feature matching is executed sequentially based on GPU-accelerated cascade hashing, in which initial candidate matches are refined by the SAO-based local constraint and RANSAC-based global verification in outlier removal. Finally, the proposed solution is evaluated and compared by using large-scale UAV datasets. The results demonstrate that it can achieve efficient feature matching with tens to hundreds of speedup ratios and provide reliable matches for SfM-based 3D reconstruction.
@@ Acknowledgments
This research was funded by the National Natural Science Foundation of China (Grant No. 42371442), Shenzhen Science and Technology Program (Grant No. JCYJ20250604181614019), and Shenzhen Key Laboratory Program (Grant No. SYSPG20241211173845013).
Appendix A
The procedure of the GPS algorithm is presented in Algorithm 2, and the code in C++ is released at https://github.com/json87/gps4mbr.
Algorithm 2 Gibbs-Poole-Stockmeyer (GPS) algorithm
1:Input: A connected graph $G=(V,E)$
2:Output: A permutation $p$ that reduces the matrix bandwidth
3:procedure PseudoPeripheralNodeSelection($G$)
4: Select an initial node $u$ with minimum degree
5: Generate the level structure $L(u)$ rooted at $u$
6:while true do
7:  Select a node $v$ in the last level of $L(u)$ with minimum degree
8:  Generate the level structure $L(v)$ rooted at $v$
9:  if depth($L(v)$) >> depth($L(u)$) then
10:   $u\leftarrow v$, $L(u)\leftarrow L(v)$
11:  else
12:   return $(u,v)$ as the pseudo-peripheral pair
13:  end if
14:end while
15:procedure GpsOrdering($G,u,v$)
16: Let $L_{u}$ and $L_{v}$ be the level structures rooted at $u$ and $v$
17: Combine $L_{u}$ and $L_{v}$ to create a combined level structure $C$
18:for each level $i$ in $C$ do
19:  Sort nodes in level $i$ based on their degrees and connection to previous levels
20:end for
21: Number nodes sequentially from level 1 to $k$ to form permutation $p$
22:return $p$